A person pulls on a container on a horizontal frictionless surface with a force of 110 N in a direction of 40° above the horizontal. If the normal force on the container is 230 N, what is the mass of the container?
Answer #1

We can resolve the force applied by person into two components.

Fx = 110 cos 40 = 84.26 N

Fy = 110 sin 40 = 70.7 N

Therefore, total normal force is

N = 70.7 + 230 = 300.7 N

we know

N = mg

m = N / g

m = 300.7 / 9.8

m = 30.68 Kg